# On the jump of the cover time in random geometric graphs

### v2

@@ 6. Upper bound right above the threshold for connectivity
It thus remains to prove Theorem 6.1. The approach of the proof is similar to the one in Section 5: first we prove that the upper bound on the effective resistance is true for any pair of vertices inside a backbone, and in a second step we then prove that the resistance between any two vertices is of the desired order. We now proceed to the details:
Proof of Theorem 6.1: Recall the definition of $\mu$ introduced in the tessellation of cubes $(\mathrm{Q_{i}})$ in Section 4 of cubes of side length $r/\mu$ (in fact cubes here are cells, but we keep the notation of the previous section). In this section, we choose $\mu(n)=\mu+O(1/n)$ to be the smallest real number larger or equal to $\mu$ such that any cube $\mathrm{Q}$ is either completely contained in $\Lambda_{n}$ or it shares at most some faces, so that we avoid boundary issues. With abuse of notation, we denote below $\mu(n)=\mu$. Let $\beta=\beta(\varepsilon)$ be a constant chosen sufficiently small (to be fixed below). We say that $\mathrm{Q_{i}}$ is dense if it has at least $\beta\log n$ vertices of $G\in\mathcal{G}(n,r,2)$, otherwise we say that it is non-dense.
See cubes $\mathrm{Q}$ as sites of $\mathbb{Z}^{2}$. The probability that a cube is non-dense is at most $n^{-\kappa}$ for some $\kappa=\kappa(\beta)>0$, since the cube $\mathrm{Q}$ has area $\Theta(\log n)$. Considering dense cubes as occupied sites of a Bernoulli site percolation model, we are thus in the super-critical phase of Bernoulli site percolation, i.e., there exists a giant component corresponding to dense cubes. Moreover, as before, it is a standard topological fact that every (maximal) connected component of cubes, which might contain dense and non-dense cubes and that has empty intersection with the giant component of dense cubes, has a diagonally connected outer boundary of non-dense cubes. As there are at most $8^{L}$ diagonally connected paths of length $L$ when the starting vertex is fixed, the probability that there exists a diagonally connected path of non-dense cubes of length greater or equal than $L$ is, by a union bound over all starting vertices, at most
$n^{1-L\beta}8^{L},$
which tends to 0 for $L=L(\beta)$ sufficiently large. Therefore, since every connected component (in the renormalized lattice of cubes $\mathrm{Q}$) of area at least $CL^{2}$ (for $C$ large enough), by a standard isoperimetric inequality has to have an outer boundary of length at least $L$, a.a.s. every component different from the giant of the associated Bernoulli site percolation has at most constant size $CL^{2}$, and Euclidean diameter at most $L$. In other words, for every site, at distance $L$ one finds a site belonging to the giant component (of the renormalized lattice), and for any vertex of the underlying random geometric graph, at Euclidean distance at most $Lr/\sqrt{5}$ we find a vertex $y$ belonging to a cube of the giant component of the Bernoulli site percolation.
Step 1: Upper bound on the effective resistance between two vertices inside the giant component of dense cubes
We first prove that the effective resistance between any two vertices inside the giant component of dense cubes is $O\left(\frac{1}{\log n}\right)$. Divide $\Lambda_{n}$ into disjoint $\alpha$-logarithmic strips of cubes $\mathrm{Q_{i}}$. We use the structure introduced in Subsection 5.1, with the difference that we consider cubes $\mathrm{Q_{i}}$ instead of cubes $\mathrm{C_{i}}$ and dense cubes playing the role of good cubes.
Let $\mathcal{Q}^{Y}_{X}$ (here $Y$ is a uniform discrete random variable independent from $X$ with values in $\{1,\dots,\lfloor c\log n\rfloor\}$, see the construction of Subsection 5.1 for the definition of $c$ and Theorem 5.1 for details of the construction of $\mathcal{Q}^{\ell}_{X}$ with a fixed value $\ell$ of $Y$) be the random path of dense cubes between two fixed dense cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$ that both belong to the giant component of dense cubes, and as in Subsection 5.1, we may assume that the picked dense cubes belong to some parallelepipeds $\mathrm{Q_{i}}\in\mathrm{T}_{s_{1},r_{1}}$ and $\mathrm{Q_{j}}\in\mathrm{T}_{s_{2},r_{2}}$ such that $r_{1}=r_{2}$, $s_{2}>s_{1}$, $s_{2}-s_{1}$ is odd, and that both parallelepipeds are not at the boundary of the 2-dimensional slice that contains them. Pick vertices $x,y$ of $G\in\mathcal{G}(n,r,2)$ such that $x\in\mathrm{Q_{i}}$ and $y\in\mathrm{Q_{j}}$. For each dense cube that is part of some realization of $\mathcal{Q}^{Y}_{X}$ select $\lfloor\beta\log n\rfloor$ vertices, except for $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$ where we may choose $\lfloor\beta\log n\rfloor-1$ vertices distinct from $x,y$ (or $\lfloor\beta\log n\rfloor-2$ in case $\mathrm{Q_{i}}=\mathrm{Q_{j}}$).
Conditional on $X$ and $Y$, we construct a random path of $G\in\mathcal{G}(n,r,2)$ that starts at $x$, finishes at $y$, and is entirely contained within the path of dense cubes $\mathcal{Q}_{X}^{Y}$ as follows. First, select uniformly at random and independently from $X,Y$ one of the $\lfloor\beta\log n\rfloor-1$ vertices that belong to the dense cube $\mathrm{Q_{i}}$ (one out of the $\lfloor\beta\log n\rfloor-1$ vertices previously selected), and connect it by an edge to the vertex $x$: this is the first edge of our path. For subsequent steps we do something similar: suppose that the last vertex that we chose belongs to the dense cube $\mathrm{Q}_{s}$. We then choose uniformly and independently at random from the $\lfloor\beta\log n\rfloor$ vertices that belong to the dense cube that follows in the path $\mathcal{L}^{Y}_{X}$. We proceed until the last selected vertex belongs to the cube just before $\mathrm{Q_{j}}$ (the dense cube that contains $y$). In this case we select uniformly at random between the $\lfloor\beta\log n\rfloor-1$ selected vertices in $\mathrm{Q_{j}}$ and finally connect all these last vertices by an edge to $y$ (if we had $\mathrm{Q_{i}}=\mathrm{Q_{j}}$, then we would pick $\lfloor\beta\log n\rfloor-2$ vertices inside the cube and form paths of length $2$ from $x$ to $y$ through these vertices). This process gives rise to $\lfloor\beta\log n\rfloor-2$ internally vertex-disjoint random paths of $G\in\mathcal{G}(n,r,2)$ with initial vertex $x$ and final vertex $y$. We denote this random path of $G\in\mathcal{G}(n,r,2)$ by $\mathcal{J}_{X,Z}^{Y}$ and denote its law by $\mathbb{P}_{X,Y,Z}$, where $Z$ denotes the randomness that comes from the path selection conditioned on $X,Y$. Observe that the number of edges that have some endpoint inside of $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$ is $\Theta(\log^{3}n)$, since $\mathrm{V}^{\prime}(\mathrm{Q_{i}})$ (see also Subsection  5.1) is made of $O(\log n)$ dense cubes, each of which has $O(\log n)$ selected vertices and between any two of such dense cubes there are $O(\log^{2}n)$ edges. Since between any two consecutive dense cubes there are a.a.s. $\Theta(\log^{2}n)$ edges, and we choose uniformly at random one of them, we obtain in the case $u\neq x,u\neq y,v\neq x,v\neq y$,
$\mathbb{P}_{X,Y,Z}(\{u,v\}\in\mathcal{J}_{Z}^{Y}|X,Y)=O\left(\frac{1}{\log^{2}n}\right),$
whereas in the case |$|\{u,v\}\cap\{x,y\}|=1$, we have
$\mathbb{P}_{X,Y,Z}(\{u,v\}\in\mathcal{J}_{X,Z}^{Y}|X,Y)=O\left(\frac{1}{\log n}\right).$
$+\sum_{t}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\end{subarray}}p^{2}(u,v).$ (6.2)
To bound the first two sums, since $p(x,v)=O\left(\frac{1}{\log n}\right)$, and since $x$ and $y$ send flow to different $O(\log n)$ vertices, we have $\sum_{v\in L_{1}}\big{(}p^{2}(y,v)+p^{2}(x,v)\big{)}=O\left(1/\log n\right)$. For the sum in the second row, since the number of edges with at least one endpoint inside of $\mathrm{V}^{\prime}(\mathrm{Q_{i}})\cup\mathrm{V}^{\prime}(\mathrm{Q_{j}})$ is $O(\log^{3}n)$, and since $p(u,v)=O(1/\log^{2}n)$,
$\sum_{\begin{subarray}{c}u\in\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\\ u\neq x,u\neq y\\ \end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ v\neq x,v\neq y\\ u\sim v\end{subarray}}p^{2}(u,v)=O\left(\frac{1}{\log n}\right).$
To bound the term in the last row, recall that, as in the proof of Theorem 5.1 in Section 5, the probability that a dense cube inside of some parallelepiped $\mathrm{T}\in D_{t}$ belongs to $\mathcal{Q}_{X}^{Y}$ is $O\left(\frac{1}{t\log n}\right)$ for $1\leq t\leq\frac{s_{2}-s_{1}+1}{2}$ and $O\left(\frac{1}{\log n}\right)$ for $t=0$, since there are $\Omega(t)$ parallelepipeds at level $t$, and we choose one of the $\Omega(\log n)$ crossings of dense cubes, uniformly at random. Now, observe that we have the additional randomness coming from the fact for any pair of consecutive dense cubes, we choose u.a.r. a vertex inside each cube, giving another factor $O(1/\log^{2}n)$, as mentioned above. All three choices are independent of each other, and thus, for edges $\{u,v\}$ appearing in the last term, we have
$p(u,v)=O\left(\frac{1}{t\log^{3}n}\right).$
The number of edges with positive probability $p$ and with at least one endpoint inside of a parallelepiped $\mathrm{T}$ is a.a.s. $O(\log^{4}n)$, as there are $O(\log^{2}n)$ dense cubes inside $\mathrm{T}$ (the volume of $\mathrm{T}$ is $\Theta(\log^{3}n)$, and each cube has area $\Theta(\log n)$), each dense cube contains $O(\log n)$ vertices, and the selected vertices are connected by an edge to another $O(\log n)$ selected vertices by construction. Thus
$\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\end{subarray}}p^{2}(u,v)$ $\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\sum_{\begin{subarray}{c}u\in\mathrm{T}_{i,j}\\ u\notin\mathrm{V}^{\prime}(\mathrm{Q}_{i})\cup\mathrm{V}^{\prime}(\mathrm{Q}_{j})\end{subarray}}\sum_{\begin{subarray}{c}v\in L_{1}\\ u\sim v\end{subarray}}\frac{c^{\prime\prime\prime}}{t^{2}\log^{6}n}$ (6.3)
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}\sum_{\mathrm{T}_{i,j}\in D_{t}}\log^{4}n\frac{c^{\prime\prime\prime\prime}}{t^{2}\log^{6}n}$
$\leq\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}(t+1)\frac{c^{\prime\prime\prime\prime}}{t^{2}\log^{2}n}$
$=O\left(\frac{1}{\log^{2}n}\right)\sum_{1\leq t\leq\frac{s_{2}-s_{1}+1}{2}}O\left(\frac{1}{t}\right)$
$=O\left(\frac{1}{\log n}\right).$
it is clear that the contribution of $D_{0}$ to the formula is $O\left(\frac{1}{\log^{2}n}\right)$. By symmetry, the contribution of $D_{t}$ with $\frac{s_{2}-s_{1}}{2}<t\leq s_{2}-s_{1}+1$ is also $O\left(\frac{1}{t\log^{2}n}\right)$, and thus
$\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{1}{\log n}\right).$
Recall that we assumed that $x,y$ belong to some dense cubes $\mathrm{Q_{i}}$ and $\mathrm{Q_{j}}$, that at the same time are inside of some parallelepipeds in the same horizontal strip and separated by an even number of parallelepipeds. Before we lift these conditions, we first provide a useful construction: we claim that if $x^{\prime},y^{\prime}\in L_{1}(G)$ with $x^{\prime}\in\mathrm{Q_{i}}^{\prime}$ and $y^{\prime}\in\mathrm{Q_{j}}^{\prime}$ with both cubes being dense, and such that $\mathrm{Q_{i}}^{\prime}$ and $\mathrm{Q_{j}}^{\prime}$ are connected by a path $\mathcal{L}$ of dense cubes of length $O(\log n)$, then $\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=\left(\frac{1}{\log n}\right)$.
To see this, let $x^{\prime}$ send an amount of $\frac{1}{\lfloor\beta\log n\rfloor-1}$ flow to the $\lfloor\beta\log n\rfloor-1$ previously selected vertices of $\mathrm{Q_{i}}^{\prime}$. From there on, let the previously selected vertices to send a uniformly distributed flow to the selected $\lfloor\beta\log n\rfloor$ vertices that belong to the dense cube that follows in the path that connects $\mathrm{Q_{i}}^{\prime}$ and $\mathrm{Q_{j}}^{\prime}$. We continue distributing in a uniform way the flow until we arrive until $\mathrm{Q_{j}}^{\prime}$. Once the flow reached the $\lfloor\beta\log n\rfloor-1$ selected vertices of $\mathrm{Q_{j}}^{\prime}$, send all the flow back to $y^{\prime}$. Denote this flow by $\theta$. We then have
As before, the first two sums contribute $O(\log n)O(1/\log^{2}n)=O(1/\log n)$. For the third sum, note that there are $O(\log n)$ dense cubes in $\mathcal{L}$, and between two such cubes there are $O(\log^{2}n)$ edges with flow $O\left(\frac{1}{\log^{2}n}\right)$ on each of them. Thus
$\mathcal{R}(x^{\prime}\longleftrightarrow y^{\prime})=O\left(\frac{1}{\log n}\right)+\sum_{|\mathcal{L}|}O\left(\log^{2}n\right)O\left(\frac{1}{\log^{4}n}\right)=O\left(\frac{1}{\log n}\right),$
and the claim is proved.
We now show how this claim allows us to prove the bound on the effective resistance for any $x\in\mathrm{C_{i}}^{\prime}\in\mathrm{T}_{s_{1},r_{1}}$ and $y\in\mathrm{C_{j}}^{\prime}\in\mathrm{T}_{s_{2},r_{2}}$ in the same strip, first assuming that both parallelepipeds do not belong to the boundary. First, if $x$ and $y$ are separated by an odd number of parallelepipeds (that is, $s_{2}-s_{1}$ even), we find $x^{\prime}$ in a parallelepiped adjacent to $x$, at Euclidean distance (in the renormalized lattice of cubes) $O(\log n)$ from $x$. Thus, by the claim, $\mathcal{R}(x\longleftrightarrow x^{\prime})=O\left(\frac{1}{\log n}\right)$, and since $x^{\prime}$ and $y$ are separated by an even number of parallelepipeds, by the above, $\mathcal{R}(x^{\prime}\longleftrightarrow y)=O\left(\frac{1}{\log n}\right)$, and finally by Lemma 2.4, also $\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{1}{\log n}\right)$. As before, by the same idea, we may extend our bounds also to arbitrary vertices inside the same parallelepiped (in cubes not satisfying all restrictions) and to vertices in parallelepipeds of a strip on the boundary.
To generalize the inequality between any pair of strips we use the same idea as in Section 5 from Figure 12: for any pair of strips, we find another vertex $z$ in the intersection of the strips that contain $x$ and $y$ and that intersect at some parallelepiped $\mathrm{T}$ that contains the vertex $z$ such that $\mathcal{R}(x\longleftrightarrow z)$ and $\mathcal{R}(y\longleftrightarrow z)$ are $O(\frac{1}{\log n})$ (such $z$ must exist, as there is a logarithmic amount of crossings of dense cubes in each strip). Together with the triangle inequality (Lemma 2.4) we thus finished proving that a.a.s. the effective resistance between any two vertices that belong to dense cubes that belong to the giant component of dense cubes is $O\left(\frac{1}{\log n}\right)$.
Step 2: Upper bound for the effective resistance between a vertex not in the giant and another in the giant
@@
Now we prove that the effective resistance between any vertex $x\in G$ that is in a cube, being dense or non-dense, that is not in the giant component of the auxiliary Bernoulli site process, and some vertex $y\in G$, belonging to a dense cube in the giant component of the auxiliary Bernoulli site process is at most $O\left(\frac{1}{\log n}\right)$: by the triangle inequality (Lemma 2.4) together with Step 1, this will then show that for any pair of vertices $x,y\in G$, we have $\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{1}{\log n}\right)$.
@@
In order to show this bound on the effective resistance, it suffices to show that for every vertex $x$ we find $y$ belonging to the giant of the auxiliary Bernoulli site process together with a set $\mathcal{P}_{x,y}$ of $\Omega(\log n)$ internal vertex-disjoint paths between $x$ and $y$ of constant length. Indeed, in this case one can split evenly the unit flow emanating from $x$ between all these paths, such that the flow on each such edge of each path is $O(1/\log n)$, and the resistance, by Definition 2.1 thus satisfies
@@
We will follow this plan in detail now. We first assume that $x$ is at distance at least $3C\sqrt{\log n}$ from all boundaries of $\Lambda_{n}$, where $C$ is a sufficiently large constant. Recall that by the beginning of the proof, for every $x\in G$ there exists $y\in G$ at distance $rL/\sqrt{5}=O(\log n)$, with $y$ belonging to a dense cube of the giant component of the auxiliary Bernoulli process (and therefore also $y\in L_{1}(G)$). We thus may find, an axis-parallel square $S$ of side length $3C\sqrt{\log n}$ containing both $x$ and $y$, such that both $x$ and $y$ are at distance at least say $C\sqrt{\log n}$ from the boundary of $S$, and such that the left bottom corner of $S$ is an integer coordinate. We now show that we cannot separate $x$ from $y$ by removing less than $\varepsilon^{\prime}\log n$ many vertices inside $S$ (and therefore we cannot separate $x$ and $y$ in $G$ either), for some $0<\varepsilon^{\prime}<\varepsilon$. Since this will hold a.a.s. for all $x,y$ simultaneously (with $S$ changing according to positions of $x,y$), we conclude that the graph is $\varepsilon^{\prime}\log n$-connected. By Menger’s theorem, for every pair of vertices there exist $\varepsilon^{\prime}\log n$ internally vertex-disjoint paths inside $S$ between any pair of vertices, and we will show then that they can be chosen of constant length.
@@ 7. Concluding remarks and future work
We showed that the cover time undergoes a jump at the threshold of connectivity, decreasing from $\Theta(n\log^{2}n)$ to $\Theta(n\log n)$ a.a.s. It is therefore natural to ask what happens right at the threshold of connectivity: assume then that $r^{d}=(\log n+f(n))/V_{d}$, where $f(n)$ may be positive or negative, and $|f(n)|=o(\log n)$. If $f(n)$ is negative and $|f(n)|=\omega(1)$, the same argument as in Case 2 of Section 4, can be used to show that the number of vertices of degree $1$ inside the giant component is a.a.s. $\Omega(nr^{d}V_{d}e^{-r^{d}V_{d}})=\Omega(e^{-f(n)})$ (which is positive since $f(n)$ is negative). Hence, we obtain a set $V^{\prime}$ of those vertices of degree 1 vertices $x,y$ for which $\log(|V^{\prime}|)=-f(n)$, and any pair of vertices of $x,y\in V^{\prime}$ satisfies $\mathcal{R}(x\longleftrightarrow y)\geq 1$. Thus, by Theorem 2.7, a.a.s. $\tau_{{\rm cov}}\geq\gamma n\log n|f(n)|$. We do not know for which values of $f(n)$ this lower bound is tight: for some values of $f(n)$ a corresponding upper bound could perhaps be obtained by splitting the cubes into certain good cubes where the flow can be nicely dissipated and bad cubes where not, and combining them in an optimal way. We do not know what happens if $f(n)$ is positive.

### v3

@@ 7. Upper bound right above the threshold for connectivity
It thus remains to prove Theorem 7.1. The approach of the proof is similar to the one in Section 6: first we prove that the upper bound on the effective resistance is true for any pair of vertices inside the backbone of dense cubes, and in a second step we then prove that the resistance between any two vertices is of the desired order. We now proceed to the details:
@@
Proof of Theorem 7.1: By Theorem 6.1, the result holds for any $x,y$ with $x\in\mathrm{Q_{i}},y\in\mathrm{Q_{j}}$, and $\mathrm{Q_{i}},\mathrm{Q_{j}}\in\mathcal{C}^{site}_{\mathrm{Q}}(n)$. It thus remains to show that the effective resistance between any vertex $x\in G$ that is in a cube, being dense or non-dense, that is not in the giant component of the auxiliary Bernoulli site process, and some vertex $y\in G$, belonging to a dense cube in the giant component of the auxiliary Bernoulli site process is at most $O\left(\frac{1}{\log n}\right)$: by the triangle inequality (Lemma 2.4) together with Theorem 6.1, this will then show that for any pair of vertices $x,y\in G$, we have $\mathcal{R}(x\longleftrightarrow y)=O\left(\frac{1}{\log n}\right)$.
@@
In order to show this bound on the effective resistance, it suffices to show that for every vertex $x$ we find $y$ belonging to the giant of the auxiliary Bernoulli site process together with a set $\mathcal{P}_{x,y}$ of $\Omega(\log n)$ internal vertex-disjoint paths between $x$ and $y$ of constant length. Indeed, in this case one can split evenly the unit flow emanating from $x$ between all these paths, such that the flow on each such edge of each path is $O(1/\log n)$, and the resistance, by Definition 2.1, thus satisfies
@@
We will follow this plan in detail now. We first assume that $x$ is at distance at least $3C\sqrt{\log n}$ from all boundaries of $\Lambda_{n}$, where $C$ is a sufficiently large constant.
Recall that by the proof of Theorem 6.1, for every $x\in G$ there exists $y\in G$ at distance $rL/\sqrt{5}=O(\log n)$, with $y$ belonging to a dense cube of the giant component of the auxiliary Bernoulli process (and therefore also $y\in L_{1}(G)$). We thus may find, an axis-parallel square $S$ of side length $3C\sqrt{\log n}$ containing both $x$ and $y$, such that both $x$ and $y$ are at distance at least say $C\sqrt{\log n}$ from the boundary of $S$, and such that the left bottom corner of $S$ is an integer coordinate. We now show that we cannot separate $x$ from $y$ by removing less than $\varepsilon^{\prime}\log n$ many vertices inside $S$ (and therefore we cannot separate $x$ and $y$ in $G$ either), for some $0<\varepsilon^{\prime}<\varepsilon$. Since this will hold a.a.s. for all $x,y$ simultaneously (with $S$ changing according to positions of $x,y$), we conclude that the graph is $\varepsilon^{\prime}\log n$-connected. By Menger’s theorem, for every pair of vertices there exist $\varepsilon^{\prime}\log n$ internally vertex-disjoint paths inside $S$ between any pair of vertices, and we will show then that they can be chosen of constant length.
@@ 8. Concluding remarks and future work
We showed that the cover time undergoes a jump at the threshold of connectivity, decreasing from $\Theta(n\log^{2}n)$ to $\Theta(n\log n)$ a.a.s. It is therefore natural to ask what happens right at the threshold of connectivity: assume then that $r^{d}=(\log n+f(n))/V_{d}$, where $f(n)$ may be positive or negative, and $|f(n)|=o(\log n)$. If $f(n)$ is negative and $|f(n)|=\omega(1)$, the same argument as in Case 2 of Section 4, can be used to show that the number of vertices of degree $1$ inside the giant component is a.a.s. $\Omega(nr^{d}V_{d}e^{-r^{d}V_{d}})=\Omega(e^{-f(n)})$ (which is positive since $f(n)$ is negative). Hence, we obtain a set $V^{\prime}$ of those vertices of degree 1 vertices $x,y$ for which $\log(|V^{\prime}|)=-f(n)$, and any pair of vertices of $x,y\in V^{\prime}$ satisfies $\mathcal{R}(x\longleftrightarrow y)\geq 1$. Thus, by Theorem 2.7, a.a.s. $\tau_{{\rm cov}}\geq\gamma n\log n|f(n)|$. We do not know for which values of $f(n)$ this lower bound is tight. On the other hand, the upper bound of Theorem 6.1 shows that between any two vertices $x,y$ belonging to a dense cube of the auxiliary giant component (as defined there), $\mathcal{R}(x\longleftrightarrow y)=O(1/\log n)$. The value of the cover time is thus determined by the maximal resistance between pairs of vertices, one of which being outside the giant. It is subject of further work to obtain results on these pairs.